what is the measure of an angle if the measure of its complement is 47 degrees

Answer by drj(1380)

What is the measure of an angle if the measure of its complement is 47 degrees

Step 1. Complementary angles means two angles add up to 90 degrees.

Step 2. Let x be the angle and 90-x be the complementary angles.

Step 3. As given, 90-x=47 or x=90-47=43

Step 4. ANSWER. Angle is 43 degrees.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

Good luck in your studies!

Respectfully,
Dr J